# Errors-In-Variables Model Fitting for Partially Unpaired Data Utilizing Mixture Models

### v2

@@ Abstract
The goal of this paper is to introduce a general argumentation framework for regression in the errors-in-variables regime, allowing for full flexibility about the dimensionality of the data, error probability density types, the (linear or nonlinear) model type and the avoidance of ad-hoc definitions of loss functions. In this framework we introduce model fitting for partially unpaired data, i.e. for given data groups the pairing information of input and output is lost (semi-supervised). This is achieved by constructing mixture model densities, which directly model this loss of pairing information allowing for inference. In a numerical simulation study linear and nonlinear model fits are illustrated as well as a real data study is presented based on life expectancy data from the world bank utilizing a multiple linear regression model. These results allow the conclusion that high quality model fitting is possible with partially unpaired data, which opens the possibility for new applications with unfortunate or deliberate loss of pairing information in the data.
@@
Dr. Högele is a Professor of Applied Mathematics and Computational Science at the Department of Computer Science and Mathematics at the Munich University of Applied Sciences HM, Germany.
Dr. Brockhaus is a Professor of Applied Mathematics and Statistics at the Department of Computer Science and Mathematics at the Munich University of Applied Sciences HM, Germany.
@@ 1 Introduction
A possibility to avoid the definition of an ad-hoc loss function is the modeling of known uncertainties in the data and applying Maximum Likelihood (ML) approaches, which, in consequence, lead inherently to data driven loss functions. A well-known connection is between a normal distributed error in the output and the squared loss function. These approaches can always be extended to Bayesian estimations if prior distributions are assumed for the model parameters leading to Maximum A Posteriori (MAP) or Minimum Mean Squared Error (MMSE) algorithms based on the posterior distribution [Hoegele et al., 2013].
@@
Unpaired / unlabeled data (also broken sampling in regression) can occur in many applications and is focus of current research, e.g. [Bai and Hsing, 2005, Liang et al., 2007, Wang et al., 2022b]. In this work, we understand by partially unpaired data (under the term semi-supervised data) specifically that for parts of the data the one-to-one pairing of input and output data is missing, but we still have paired subgroups in the data. Semi-supervised data often is understood as having additional input data without output data, e.g. [Kostopoulos et al., 2018, Qi and Luo, 2022]. This is different to the partially unpaired data in this paper and we regard this as a different flavor of semi-supervised data since we lose significant information compared to a fully supervised framework, but we still have (possibly weak) input/output relations in contrast to unsupervised learning. See Figure 1 for an illustration. It is demonstrated that mixtures of labeled and unlabeled data can improve the predictive performance in regression problems [Liang et al., 2007]. There are different strategies to deal with such deficiencies, such as altering the data set by data imputation approaches, e.g. see [Bennett, 2001, Sterne et al., 2009], or focusing on unordered subsets in hypothesis testing [Wang et al., 2022b]. We regard the work of Liang et al. [Liang et al., 2007] considering a general predictive Bayesian frameworks for mixed labeled and unlabeled data as closest to our goal. Although the work provides a very general framework for regression and classification, it misses the configurational complexity of partially unpaired data. In consequence, one task of this paper is to incorporate the largest possible variety of missing pairing information transparently to model fitting by avoiding data alteration (deletion or imputation) and actively constructing mixture model probability densities accurately representing the partial pairing.
@@
Mixture Models are applied in a variety of applications, especially Gaussian Mixture Models (GMM) are very popular, e.g. [Liang et al., 2007]. A main task is in the literature to find the GMM description of a given data set utilizing expectation maximization (EM) algorithms, also for the task of missing data deficiencies [Michael et al., 2020, McCaw et al., 2022]. To the knowledge of the authors, a general framework for constructing mixture models for dealing with the high complexity of corrupt / incomplete pairing information of the data with the goal to support a model fitting problem is missing in the literature. Closely related approaches of utilizing mixture models in applied mathematics with lost pairing information are presented for computer vision in order to deal with unknown feature correspondences [Hoegele, 2024a] and for random equations with high combinatorial possibilities for the stochastic parameters [Hoegele, 2024b].
@@
In this paper, we propose to consider the problem of model fitting in a new conclusive way. A general probabilistic framework for fitting models in data based purely on observational probability density functions including errors-in-variables is presented, which has direct relations to well-known standard methods for completely paired data, such as ordinary least squares, Deming regression [Deming, 1964], interval data regression as well as multiple linear regression. This framework will be generalized from supervised to semi-supervised model fitting by including (partially) unpaired data in one common line of stochastic argumentation. The derivations allow for flexibility about i) the number and dimensions of the input / output data, ii) the type of individual error characteristics of each data point with errors-in-variables utilizing general density functions, iii) the type of (linear or nonlinear) models which should be fitted and iv) the pairing information level. In summary, there are two main goals of this paper: A) presenting a general stochastic argumentation framework for model fitting (without ad-hoc loss functions) and B) presenting an extension of model fitting within this framework to partially unpaired data utilizing mixture models.
@@
In A) the main ideas of the stochastic argumentation are: i) Formulating the fitting problem in equations containing all random variables, which equals to $\boldsymbol{0}$, and identifying this as a Maximum Likelihood (ML) problem. ii) Applying the law of total probability for densities wherever necessary to make sure that correct stochastic dependencies are utilized and identifying the density functions of the basic random variables in the ML problem. These two ideas are presented repeatedly for different fitting scenarios and their generality is a first main result of this paper.
@@
In B) the extension of model fitting to partially unpaired data is structured in the following three-step-approach: Presenting the cases for
@@
In the results section, we demonstrate the applicability of this framework by simulation studies for a line fit (Section 3.1) and a fit of anisotropic noisy data with a cubic polynomial (Section 3.2). Further, in Section 4, we will demonstrate how this argumentation can be applied to multiple linear regression for life expectancy data from the world bank. These results demonstrate the importance of modeling the inherent uncertainties and the use of different levels of pairing information in data.
@@ 2.5.3 Bayesian Extension
with $c$ a normalization constant. By utilizing a non-informative prior $\pi(\boldsymbol{\alpha})$ (such as $\pi(\boldsymbol{\alpha})=$ const. on a large enough domain) the previously presented optimization problem is identical to the maximization of the posterior, leading to a Maximum A Posteriori (MAP) estimate. This allows to interpret the plotted objective functions in the results Section 3.1 as presentations of the density functions of $\boldsymbol{\alpha}$, containing directly the inherent uncertainties about $\boldsymbol{\alpha}$ graphically as intensity maps.
@@ 3.1 Demonstration for a Line Fit
Qualitative conclusions from Figures 2 and 3: First, utilizing completely unpaired data leads to arbitrary insufficient results which can be observed by the non-distinct maxima of the objective functions. Second, the fewer subgroups $R$ are utilized, the broader (and more uncertain) gets the maximum in the objective function in Figure 2. Further, for very few subgroups, such as $R=3$, a bias on the slope $\alpha_{2}$ is introduced for all scenarios as presented in Figure 3. On the other side, it is obvious that reasonable estimation of the parameters is absolutely possible even if only partially unpaired data is available (comparing $R=12$ and $R=300$). Third, comparing scenarios $A$ to $B$: The uncertainty of estimation increases with an increased noise level of the data for all cases of $R$. Fourth, comparing scenarios $B$ to $C$: The higher noise level in the data leads to similar uncertainties in comparison to fewer data. Fifth, comparing scenarios $A$ to $D$: The main difference utilizing interval data compared to Gaussian disturbances is that also for the completely paired case no distinct maximum appears but a plateau of high intensity values are observable in the objective function in Figure 2. In Figure 3 it can be seen that interval data leads to similar results with a stronger bias on the slope for $R=3$.
@@ 5 Discussion and Conclusion
In this work, we presented a general framework for model fitting scenarios with stochastic disturbances and uncertainties for completely paired and partially unpaired data utilizing mixture models. The main advantage of this approach is its generality allowing for flexibility about i) the number and dimensions of the data points $\boldsymbol{x}_{h}$ and $\boldsymbol{y}_{l}$, ii) the (possibly) individual error characteristics of each data point in the errors-in-variables framework described by the general density functions $f_{\boldsymbol{\eta}_{h}}(\boldsymbol{s})$ and $f_{\boldsymbol{\varepsilon}_{l}}(\boldsymbol{s})$, iii) the type of (linear or nonlinear) model $\boldsymbol{M}(\;\cdot\;;\boldsymbol{\alpha})$, iv) the specific form of the pairing information level and number of elements in each unpaired subgroup, and v) completely avoids ad-hoc loss functions. To be more precise: The loss function is itself derived from the structure of the pairing of the data and their (possibly) individual error characteristics and, in consequence, represents the most suitable fitting approach for such problems. This allows a very broad application of this argumentation framework and includes many of the known fitting approaches as special cases.
@@
In the paper, it is only briefly presented how the ML approach can be directly extended to Bayesian model parameter estimation utilizing Bayes’ theorem by assuming prior distributions on the parameters and applying MAP and MMSE approaches. In the context of this work, this was done in order to directly interpret the plotted likelihood functions in the results section as posterior densities by utilizing a non-informative prior. With this interpretation, we are able to directly quantify the uncertainties of the parameters $\boldsymbol{\alpha}$ of the model fit, allowing the calculation of credibility intervals or regions.

### v3

@@ Abstract
We introduce a general framework for regression in the errors-in-variables regime, allowing for full flexibility about the dimensionality of the data, observational error probability density types, the (nonlinear) model type and the avoidance of ad-hoc definitions of loss functions. In this framework, we introduce model fitting for partially unpaired data, i.e. for given data groups the pairing information of input and output is lost (semi-supervised). This is achieved by constructing mixture model densities, which directly model the loss of pairing information allowing inference. In a numerical simulation study linear and nonlinear model fits are illustrated as well as a real data study is presented based on life expectancy data from the world bank utilizing a multiple linear regression model. These results show that high quality model fitting is possible with partially unpaired data, which opens the possibility for new applications with unfortunate or deliberate loss of pairing information in data.
@@
Dr. Högele is a Full Professor of Applied Mathematics and Computational Science at the Department of Computer Science and Mathematics at the Munich University of Applied Sciences HM, Germany.
Dr. Brockhaus is a Full Professor of Applied Mathematics and Statistics at the Department of Computer Science and Mathematics at the Munich University of Applied Sciences HM, Germany.
The authors report there are no competing interests to declare.
This manuscript appears on ArXiv.org
arxiv:2406.18154 [stat.ME], https://doi.org/10.48550/arXiv.2406.18154
@@ 1 Introduction
A possibility to avoid the definition of an ad-hoc loss function is the modeling of known uncertainties in the data and applying Maximum Likelihood (ML) approaches, which, in consequence, lead inherently to data driven loss functions. For example, data uncertainties can appear only in the output data (such as for ordinary least squares) or in input and output data, also known as the errors-in-variables approaches. A well-known connection is between a normal distributed error in the output and the squared loss function. These approaches can always be extended to Bayesian estimations if prior distributions are assumed for the model parameters leading to Maximum A Posteriori (MAP) or Minimum Mean Squared Error (MMSE) algorithms based on the posterior distribution [Hoegele et al., 2013].
@@
Unpaired / unlabeled data (also broken sampling in regression) can occur in many applications and is focus of current research, e.g. [Bai and Hsing, 2005, Liang et al., 2007, Wang et al., 2022b]. In this work, we understand by partially unpaired data (under the term semi-supervised data) specifically that for parts of the data the one-to-one pairing of input and output data is missing, but we still have paired subgroups in the data. Semi-supervised data often is understood as having additional input data without output data, e.g. [Kostopoulos et al., 2018, Qi and Luo, 2022]. This is different to the partially unpaired data in this paper and we regard this as a different flavor of semi-supervised data since we lose significant information compared to a fully supervised framework, but we still have (possibly weak) input/output relations in contrast to unsupervised learning. See Figure 1 for an illustration of different pairing information between input $\boldsymbol{x}$ and output $\boldsymbol{y}$. It is demonstrated that mixtures of labeled and unlabeled data can improve the predictive performance in regression problems [Liang et al., 2007]. There are different strategies to deal with such deficiencies, such as altering the data set by data imputation approaches, e.g. see [Bennett, 2001, Sterne et al., 2009], or focusing on unordered subsets in hypothesis testing [Wang et al., 2022b]. We regard the work of Liang et al. [Liang et al., 2007] considering a general predictive Bayesian frameworks for mixed labeled and unlabeled data as closest to our goal. Although the work provides a very general framework for regression and classification, it misses the configurational complexity of partially unpaired data. In consequence, one task of this paper is to incorporate the largest possible variety of missing pairing information transparently to model fitting by avoiding data alteration (deletion or imputation) and actively constructing mixture model probability densities accurately representing the partial pairing.
@@
Mixture Models are probability density functions, which are composed of the weighted sum of elementary probability densities. They are applied in a variety of applications, especially Gaussian Mixture Models (GMM) are very popular, which are the weighted sum of Gaussian densities, e.g. [Liang et al., 2007]. A main task is in the literature to find the GMM description of a given data set utilizing expectation maximization (EM) algorithms, also for the task of missing data deficiencies [Michael et al., 2020, McCaw et al., 2022]. To the knowledge of the authors, a general framework for constructing mixture models for dealing with the high complexity of corrupt / incomplete pairing information of the data with the goal to support a model fitting problem is missing in the literature. Closely related approaches of utilizing mixture models in applied mathematics with lost pairing information are presented for computer vision in order to deal with unknown feature correspondences [Hoegele, 2024a] and for random equations with high combinatorial possibilities for the stochastic parameters [Hoegele, 2024b].
@@
In this paper, we propose to consider the problem of model fitting in a new conclusive way. A general probabilistic framework for fitting models in data based purely on observational error probability density functions including errors-in-variables is presented, which has direct relations to well-known standard methods for completely paired data, such as ordinary least squares, Deming regression [Deming, 1964], total least squares[Markovsky and Van Huffel, 2007], interval data regression as well as multiple linear regression. Observational error densities can have different reasons, for example, errors only in the output can be classical Gaussian measurement errors, but if regression is performed additionally with measured input data then both, input and output errors (errors-in-variables), are typically described by Gaussians. Another example is interval data, which can be represented by uniform densitiy functions and which can occur, e.g., in survey data, in particular, when asking for sensitive information like income. This framework will be generalized from supervised to semi-supervised model fitting by including (partially) unpaired data in one common line of stochastic argumentation. It is a key point that the presented derivations allow for full flexibility about i) the number and dimensions of the input / output data, ii) the type of individual error characteristics of each data point with errors-in-variables utilizing general density functions, iii) the type of (linear or nonlinear) models which should be fitted and iv) the pairing information level. In the schematic Figure 2 the concepts of this work are explained in an overview starting from the well–known ordinary least squares application of a line fit in Subfigure A to the most general concept of this paper in Subfigure D with a nonlinear fit in partially unpaired data. This presentation utilizes a one-dimensional input and output for illustrative purposes, which is generalized in the paper to arbitrary dimensions.
Figure 2: Schematic illustration of the model fitting concepts in this paper for an input $x\in\mathbb{R}$ and output $y\in\mathbb{R}$, presented as steps of generalizations. A) Well–known ordinary least squares for a line fit model, which corresponds to Maximum Likelihood estimation with a constant Gaussian observational error density $f_{\boldsymbol{\varepsilon}}$ in $y$. B) Generalization of case (A) to a general nonlinear model $M(x;\boldsymbol{\alpha})$ and general observational error density $f_{\boldsymbol{\varepsilon}_{l}}$ in $y$ individual for each data point $l=1,..,L$. This is presented in Supplement A. C) Generalization of (B) to the errors-in-variables approach including observational error density functions $f_{\boldsymbol{\eta}_{l}}$ in $x$. This is presented in Section 2.2. D) Generalization of (C) to also include partially unpaired data subgroups, which contains all possible pairings of data within a subgroup (in this case $3\cdot 3=9$ pairings). This is realized with mixture models, which are the weighted sum of the data point probability densities. This is presented in Section 2.4.
In summary, there are two main goals of this paper: I) presenting a general stochastic argumentation framework for model fitting (without ad-hoc loss functions) (see Figure 2A–C) and II) presenting an extension of model fitting within this framework to partially unpaired data utilizing mixture models (see Figure 2D).
@@
In I) the main ideas of the stochastic argumentation are: a) Formulating the fitting problem in equations containing all random variables, which equal to $\boldsymbol{0}$, and identifying this as a Maximum Likelihood (ML) problem. b) Applying the law of total probability for densities wherever necessary to make sure that correct stochastic dependencies are utilized and identifying the density functions of the basic random variables in the ML problem. c) Presentation of the optimization problems by the resulting objective functions. Points a) to c) are presented repeatedly for different fitting scenarios and their generality is a first main result of this paper.
@@
In II) the extension of model fitting to partially unpaired data is structured in the following three-step-approach: Presenting the cases for
@@
In the results section, we demonstrate the applicability of this framework by simulation studies with Gaussian and uniform mixture models as observational error densities for a line fit (Section 3.1) and a fit of anisotropic noisy data with a cubic polynomial (Section 3.2). Further, in Section 4, we will demonstrate how this argumentation can be applied to multiple linear regression for life expectancy data from the world bank. These results demonstrate the importance of modeling the inherent uncertainties and the use of different levels of pairing information in data.
@@ 2.5.3 Bayesian Extension
with $c$ a normalization constant. By utilizing a non- or weakly informative prior $\pi(\boldsymbol{\alpha})$ (such as $\pi(\boldsymbol{\alpha})=$ const. on a large enough domain) the previously presented optimization problem is identical to the maximization of the posterior, leading to a Maximum A Posteriori (MAP) estimate. This allows to interpret the plotted objective functions in the results Section 3.1 as presentations of the density functions of $\boldsymbol{\alpha}$, containing directly the inherent uncertainties about $\boldsymbol{\alpha}$ graphically as intensity maps.
@@ 3.1 Demonstration for a Line Fit
Qualitative conclusions from Figures 3 and 4: First, utilizing completely unpaired data leads to arbitrary insufficient results which can be observed by the non-distinct maxima of the objective functions. Second, the fewer subgroups $R$ are utilized, the broader (and more uncertain) gets the maximum in the objective function in Figure 3. Further, for very few subgroups, such as $R=3$, a bias on the slope $\alpha_{2}$ is introduced for all scenarios as presented in Figure 4. On the other side, it is obvious that reasonable estimation of the parameters is absolutely possible even if only partially unpaired data is available (comparing $R=12$ and $R=300$). Third, comparing scenarios $A$ to $B$: The uncertainty of estimation increases with an increased noise level of the data for all cases of $R$. Fourth, comparing scenarios $B$ to $C$: The higher noise level in the data leads to similar uncertainties in comparison to fewer data. Fifth, comparing scenarios $A$ to $D$: The main difference utilizing interval data compared to Gaussian disturbances is that also for the completely paired case no distinct maximum appears but a plateau of high intensity values are observable in the objective function in Figure 3. In Figure 4 it can be seen that interval data leads to similar results with a stronger bias on the slope for $R=3$.
@@ 5 Discussion and Conclusion
In this work, we presented a general framework for model fitting scenarios with stochastic disturbances and uncertainties for completely paired and partially unpaired data utilizing mixture models. The main advantage of this approach is its generality allowing for full flexibility about i) the number and dimensions of the data points $\boldsymbol{x}_{h}$ and $\boldsymbol{y}_{l}$, ii) the (possibly) individual error characteristics of each data point in the errors-in-variables framework described by the general density functions $f_{\boldsymbol{\eta}_{h}}(\boldsymbol{s})$ and $f_{\boldsymbol{\varepsilon}_{l}}(\boldsymbol{s})$, iii) the type of (linear or nonlinear) model $\boldsymbol{M}(\;\cdot\;;\boldsymbol{\alpha})$, iv) the specific form of the pairing information level and number of elements in each unpaired subgroup, and v) completely avoids ad-hoc loss functions. To be more precise: The loss function is itself derived from the structure of the pairing of the data and their (possibly) individual error characteristics and, in consequence, represents the most suitable fitting approach for such problems. This allows a very broad application of this argumentation framework and includes many of the known fitting approaches as special cases.
@@
In the paper, it is only briefly presented how the ML approach can be directly extended to Bayesian model parameter estimation utilizing Bayes’ theorem by assuming prior distributions on the parameters and applying MAP and MMSE approaches. In the context of this work, this was done in order to directly interpret the plotted likelihood functions in the results section as posterior densities by utilizing a non- or weakly informative prior. With this interpretation, we are able to directly quantify the uncertainties of the parameters $\boldsymbol{\alpha}$ of the model fit, allowing the calculation of credibility intervals or regions.